2012年2月10日星期五

10/2/2012 Week2

Today we corrected the way to test the PIR sensor. Then we checked the primary circuit and found it can work sucessfully. When the oscilloscope sets to 0.2s and 5mV, the original Vpp of PIR sensor is 10mV.

When it sets to 0.2s and 20mV, the Vpp of the primary circut is about 124mV. The votage gain is nearly equal to the theoretical value 13. And we tested the filter and found the output signal was clearer than before. However, the second part of the circuit still not work.

Then we modified the second part by using another new design as below, but it still does not work.



After that, we changed the experimental position, because we want to store images by digital oscilloscope. Strangely,the values displayed on this oscilloscope are different from former one.





We decided to write the report first and test the circuit next time.

没有评论:

发表评论